# MCbiF: Measuring Topological Autocorrelation in Multiscale Clusterings via 2-Parameter Persistent Homology

### v1

@@ 2. C2:
Next, we provide the proof of Proposition 17 about the relation between triangle 0-conflicts and the clustering coefficient.
@@
Assume that $\mathcal{C}(K_{1}^{s,t})<1$. Then there exist $x,y,z\in X$ that form a path of length 2 but no triangle, see Newman [2018] for details on the clustering coefficient. Without loss of generality, $x,y],[y,z]\in K_{1}^{s,t}$ but $x,z]\notin K_{1}^{s,t}$. This implies $\exists r_{1},r_{2}\in[s,t]$: $x\sim_{r_{1}}y\sim_{r_{2}}z$ and $\nexists r\in[s,t]$: $x\sim_{r}z$. Hence, $x,y,z$ lead to a triangle 0-conflict. ∎
We can now prove Proposition 19 on 1-conflicts.
@@
We next prove statement (ii): If $\mathrm{HF}_{1}(s,t)\geq 1$ there exists a 1-cycle $z=[x_{1},x_{2}]+\dots+[x_{n-1},x_{n}]+[x_{n},x_{1}]$ that is non-bounding, i.e., $h:=[z]\neq 0$ in $H_{1}(K^{s,t})$, see Appendix C for details. Case 1: Assume $\nexists r\in[s,t]$: $x_{1}\sim_{r}x_{2}\sim_{r}x_{3}$, then it follows immediately that $x_{1},x_{2},x_{3}$ lead to a triangle 0-conflict. Case 2: Assume $\exists r\in[s,t]$: $x_{1}\sim_{r}x_{2}\sim_{r}x_{3}$. As $z]\neq 0$ there exists a 1-cycle $\tilde{z}=[\tilde{x}_{1},\tilde{x}_{2}]+\dots++[\tilde{x}_{m-1},\tilde{x}_{m}]+[\tilde{x}_{m},\tilde{x}_{1}]\in Z_{1}(K^{s,t})$ such that $\tilde{z}$ is homologous to $z$, i.e., $\tilde{z}=z+\partial_{2}w$ for $w\in C_{2}(K^{s,t})$, and such that $\nexists r\in[s,t]$: $\tilde{x}_{1}\sim_{r}\tilde{x}_{2}\sim_{r}\tilde{x}_{3}$. In particular, $\tilde{x}_{1},\tilde{x}_{2},\tilde{x}_{3}$ lead to a triangle 0-conflict.
@@
A.2.2 Proofs for Section 4.2
@@
The proof follows from Proposition 30 in Schindler and Barahona [2025], which extends directly to the 2-parameter case. ∎
@@
Statement (i) follows directly from the definition in Eq. (5). We show statement (ii) by induction. Base case: From the definition of the nerve-based MCbiF, it follows directly that $\dim N^{t_{m},t_{m}}=0$ because the indices in $A(m,m)$ correspond to mutually exclusive clusters. Induction step: Let us assume that $\dim N^{t_{m},t_{m+n}}=n$, then there exist $C_{0},\dots,C_{n}\in\theta([t_{m},t_{m+n}])$ such that $C_{0}\cap\dots\cap C_{n}\neq\emptyset$. As the clusters in partition $\theta(t_{m+n+1})$ cover the set $X$ there exist a cluster $C\in\theta(t_{m+n+1})$ such that $C\cap C_{0}\cap\dots\cap C_{n}\neq\emptyset$. Hence, $\dim N^{t_{m},t_{m}+n}\geq n+1$. If $\dim N^{t_{m},t_{m}+n}>n+1$ there would exist a second cluster $C^{\prime}\in\theta(t_{m+n+1})$ with $C^{\prime}\cap C\cap C_{0}\cap\dots\cap C_{n}\neq\emptyset$ but $C^{\prime}\cap C\neq\emptyset$ contradicts that clusters of $\theta(t_{m+n+1})$ are mutually exclusive. Hence, $\dim N^{t_{m},t_{m}+n}=n+1$, proving statement (ii) by induction. ∎
@@
The Sankey diagram graph $S(\theta)=(V=V_{1}\uplus...\uplus V_{M},E=E_{1}\uplus...\uplus E_{M-1})$ is a strict 1-dimensional subcomplex of $\tilde{K}=\tilde{K}^{t_{1},t_{M}}$ because $\tilde{K}^{t_{m},t_{m}}=V_{m}\subseteq\tilde{K}$ and $\tilde{K}^{t_{m},t_{m+1}}=E_{m}\subseteq\tilde{K}$. This also shows that the zigzag filtration (9) contains exactly the same vertices (0-simplices) and edges (1-simplices) as $S(\theta)$. ∎
@@ A.3 Proofs for Section 5
Next, we provide a proof of Proposition 32 that establishes the connection between the MPH of the MCbiF and that of the Merge-Rips bifiltration constructed from the matrix of first-merge times, $D_{\theta,s}$.
@@
The proof of the proposition then follows from a simple extension of Proposition 32 in Schindler and Barahona [2025] to the 2-parameter case. To see that the 0-dimensional MPH of $\mathcal{L}$ and $\mathcal{M}$ are equivalent, note that both bifiltrations have the same 1-skeleton. Moreover, the 1-dimensional MPH is generally not equivalent because $\mathcal{L}$ is a Rips-based bifiltration and thus 2-determined, whereas $\mathcal{M}$ is not 2-determined.
If $\theta$ is strictly hierarchical, then $D_{\theta,s}$ fulfils the strong-triangle inequality and thus the Rips-based bifiltration leads to a trivial 1-dimensional homology, see [Schindler and Barahona, 2025, Corollary 33]. Hence, the 1-dimensional MPH of $\mathcal{L}$ is trivial. ∎
@@
Using Proposition 27, we prove the statement with the equivalent nerve-based MCbiF. Note that the graph $G:=P_{t_{2}|t_{1}}P_{t_{2}|t_{1}}^{T}$ has the same vertices and edges as the simplicial complex $\tilde{K}^{t_{1},t_{2}}$, which is 1-dimensional and thus also a graph according to Proposition 26. This shows that $\mathrm{HF}_{(}t_{1},t_{2})$ is given by the number of connected components in $G$. Furthermore, observe that $P_{t_{2}|t_{1}}^{T}\mathbf{1}=\mathbf{1}$, and that the resulting matrix $L$ is the Laplacian of the undirected graph $G$. Hence, $\dim(\ker L)$ is equal to the number of connected graph components [Chung, 1997], proving the statement. ∎
A.4 Proofs for Section 7
@@
Appendix B Additional Examples
@@
Figure 9: Hilbert functions and Sankey diagram for the sequence of partitions $\theta$ defined in Example 43.
@@
Let $X=\{x_{1},x_{2},x_{3},x_{4}\}$ for which we consider two different sequences of partitions $\theta(t)$ and $\eta(t)$ such that $\theta(1)=\eta(1)=\{\{x_{1},x_{2}\},\{x_{3}\},\{x_{4}\}\}$, $\theta(2)=\eta(2)=\{\{x_{1}\},\{x_{2},x_{3}\},\{x_{4}\}\}$ but $\theta(3)=\{\{x_{1},x_{3}\},\{x_{2}\},\{x_{4}\}\}\neq\theta(3)=\{\{x_{1}\},\{x_{2}\},\{x_{3},x_{4}\}\}$. See Fig. 10 for a Sankey diagram representation of the two diagrams. Note that $\theta$ and $\eta$ only differ at scale $t=3$. However, this difference is crucial because a 1-conflict emerges in $\theta$ at scale $t=0$, whereas $\eta$ has only triangle 0-conflicts and no 1-conflict. Note that $\theta$ corresponds to the toy example in Fig. 1 with one additional isolated element.

### v2

@@ 2. C2:
We can further detect triangle 0-conflicts by analysing the graph-theoretic properties of the MCbiF 1-skeleton $K_{1}^{s,t}$. Recall that the clustering coefficient $\mathrm{C}$ of a graph is defined as the ratio of the number of triangles to the number of paths of length 2 in the graph  (Luce & Perry, 1949; Newman, 2018).
Proposition 30.
@@
Assume that $\mathcal{C}(K_{1}^{s,t})<1$. Then there exist $x,y,z\in X$ that form a path of length 2 but no triangle, see Newman (2018) for details on the clustering coefficient. Without loss of generality, $x,y],[y,z]\in K_{1}^{s,t}$ but $x,z]\notin K_{1}^{s,t}$. This implies $\exists r_{1},r_{2}\in[s,t]$: $x\sim_{r_{1}}y\sim_{r_{2}}z$ and $\nexists r\in[s,t]$: $x\sim_{r}z$. Hence, $x,y,z$ lead to a triangle 0-conflict. ∎
Let us consider the graph generated as the disjoint union of all clusters from partitions in [s,t] as cliques. This graph is equivalent to the MCbiF 1-skeleton $K_{1}^{s,t}$. Proposition 30 shows that the clustering coefficient of this graph can be used to detect triangle 0-conflicts that are not cycles. To be able to detect triangle 0-conflicts that correspond to non-bounding cycles, we turn to the 1-dimensional homology.
We next prove Proposition 11 on 1-conflicts.
@@
We next prove statement (ii): If $\mathrm{HF}_{1}(s,t)\geq 1$ there exists a 1-cycle $z=[x_{1},x_{2}]+\dots+[x_{n-1},x_{n}]+[x_{n},x_{1}]$ that is non-bounding, i.e., $h:=[z]\neq 0$ in $H_{1}(K^{s,t})$, see Appendix E.3 for details. Case 1: Assume $\nexists r\in[s,t]$: $x_{1}\sim_{r}x_{2}\sim_{r}x_{3}$, then it follows immediately that $x_{1},x_{2},x_{3}$ lead to a triangle 0-conflict. Case 2: Assume $\exists r\in[s,t]$: $x_{1}\sim_{r}x_{2}\sim_{r}x_{3}$. As $z]\neq 0$ there exists a 1-cycle $\tilde{z}=[\tilde{x}_{1},\tilde{x}_{2}]+\dots++[\tilde{x}_{m-1},\tilde{x}_{m}]+[\tilde{x}_{m},\tilde{x}_{1}]\in Z_{1}(K^{s,t})$ such that $\tilde{z}$ is homologous to $z$, i.e., $\tilde{z}=z+\partial_{2}w$ for $w\in C_{2}(K^{s,t})$, and such that $\nexists r\in[s,t]$: $\tilde{x}_{1}\sim_{r}\tilde{x}_{2}\sim_{r}\tilde{x}_{3}$. In particular, $\tilde{x}_{1},\tilde{x}_{2},\tilde{x}_{3}$ lead to a triangle 0-conflict.
@@
Remark 31.
Remark 32.
B.2.2 Proofs and Details for Section 3.2
To prove the equivalence between the MPH of the MCbiF and nerve-based MPH, we can use a Persistent Nerve Lemma for abstract simplicial complexes by Schindler & Barahona (2025).
Lemma 33 (Schindler & Barahona (2025), Lemma 32).
Let $K\subseteq K^{\prime}$ be two finite abstract simplicial complexes and $\{K_{\alpha}\}_{\alpha\in A}$ and $\{K^{\prime}_{\alpha}\}_{\alpha\in A}$ be subcomplexes that cover $K$ and $K^{\prime}$ respectively, based on the same finite parameter set such that $K_{\alpha}\subseteq K^{\prime}_{\alpha}$ for all $\alpha\in A$. Let $\tilde{K}$ denote the nerve $\mathcal{N}\left(\{|K_{\alpha}|\}_{\alpha\in A}\right)$ and $\tilde{K}^{\prime}$ the nerve $\mathcal{N}\left(\{|K^{\prime}_{\alpha}|\}_{\alpha\in A}\right)$. If the intersections $\bigcap_{i=0}^{k}|K_{\alpha_{i}}|$ and $\bigcap_{i=0}^{k}|K^{\prime}_{\alpha_{i}}|$ are either empty or contractible for all $k\in\mathbbm{N}$ and for all $\alpha_{0},...,\alpha_{k}\in A$, then there exist homotopy equivalences $\tilde{K}\rightarrow|K|$ and $\tilde{K}^{\prime}\rightarrow|K^{\prime}|$ that commute with the canonical inclusions $|K|\hookrightarrow|K^{\prime}|$ and $\tilde{K}\hookrightarrow\tilde{K}^{\prime}$.
See Schindler & Barahona (2025, Lemma 32) for a proof.
@@
We use Lemma 33 to show the equivalence between MCbiF and nerve-based MCbiF. For $s\geq s^{\prime}$ and $t\leq t^{\prime}$, we denote $\tilde{K}:=\tilde{K}^{s,t}$ and $K:=K^{s,t}$, and further denote $\tilde{K}^{\prime}:=\tilde{K}^{s^{\prime},t^{\prime}}$ and $K^{\prime}:=K^{s^{\prime},t^{\prime}}$. As $\theta$ is a piecewise-constant function with change points $t_{1}<\dots<t_{M}$, we can assume without loss of generality that $s^{\prime}=t_{k}$, $s=t_{\ell}$, $t_{m}, t=t$$^{\prime}=t_{n}$ for change points $t_{k}<t_{\ell}<t_{m}<t_{n}$ of $\theta$. For the multi-index set $A:=A(k,n)$, define the cover $\{K_{\alpha}\}_{\alpha\in A}$ by $K_{\alpha}=\Delta C_{\alpha}$ if $\alpha\in A(\ell,m)\subseteq A$ and $K_{\alpha}=\emptyset$ otherwise and the cover $\{K^{\prime}_{\alpha}\}_{\alpha\in A}$ by $K^{\prime}_{\alpha}=\Delta C_{\alpha}$ for all $\alpha\in A$. Then we have $K_{\alpha}\subseteq K^{\prime}_{\alpha}$ for all $\alpha\in A$ and we recover the MCbiF $K=\bigcup_{\alpha\in A}K_{\alpha}$ and the nerve-based MCbiF $\tilde{K}=\mathcal{N}\left(\{K^{\prime}_{\alpha}\}_{\alpha\in A}\right)$ and similarly we recover $K^{\prime}$ and $\tilde{K}^{\prime}$. For any $k\in\mathbbm{N}$ and $\alpha_{0},...,\alpha_{k}\in A$, the intersections $\bigcap_{i=0}^{k}|K^{\prime}_{\alpha_{i}}|$ are intersections of solid simplices and thus either empty or contractible. For a detailed proof of this fact, see the proof of Proposition 30 in Schindler & Barahona (2025). Lemma 33 now yields homotopy equivalences $\tilde{K}\rightarrow|K|$ and $\tilde{K}^{\prime}\rightarrow|K^{\prime}|$ that commute with the canonical inclusions $|K|\hookrightarrow|K^{\prime}|$ and $\tilde{K}\hookrightarrow\tilde{K}^{\prime}$. This shows the equivalence between the MPH of the MCbiF and the MPH of the nerve-based MCbiF. ∎
@@
Statement (i) follows directly from the definition in Eq. equation 1. We show statement (ii) by induction. Base case: From the definition of the nerve-based MCbiF, it follows directly that $\dim N^{t_{m},t_{m}}=0$ because the indices in $A(m,m)$ correspond to mutually exclusive clusters. Induction step: Let us assume that $\dim N^{t_{m},t_{m+n}}=n$, then there exist $C_{0},\dots,C_{n}\in\theta([t_{m},t_{m+n}])$ such that $C_{0}\cap\dots\cap C_{n}\neq\emptyset$. As the clusters in partition $\theta(t_{m+n+1})$ cover the set $X$ there exist a cluster $C\in\theta(t_{m+n+1})$ such that $C\cap C_{0}\cap\dots\cap C_{n}\neq\emptyset$. Hence, $\dim N^{t_{m},t_{m}+n}\geq n+1$. If $\dim N^{t_{m},t_{m}+n}>n+1$ there would exist a second cluster $C^{\prime}\in\theta(t_{m+n+1})$ with $C^{\prime}\cap C\cap C_{0}\cap\dots\cap C_{n}\neq\emptyset$ but $C^{\prime}\cap C\neq\emptyset$ contradicts that clusters of $\theta(t_{m+n+1})$ are mutually exclusive. Hence, $\dim N^{t_{m},t_{m}+n}=n+1$, proving statement (ii) by induction. ∎
@@
The Sankey diagram graph $S(\theta)=(V=V_{1}\uplus...\uplus V_{M},E=E_{1}\uplus...\uplus E_{M-1})$ is a strict 1-dimensional subcomplex of $\tilde{K}=\tilde{K}^{t_{1},t_{M}}$ because $\tilde{K}^{t_{m},t_{m}}=V_{m}\subseteq\tilde{K}$ and $\tilde{K}^{t_{m},t_{m+1}}=E_{m}\subseteq\tilde{K}$. This also shows that the zigzag filtration equation 5 contains exactly the same vertices (0-simplices) and edges (1-simplices) as $S(\theta)$. ∎
@@ B.3 Proofs and Details for Section 4
Next, we state Proposition 35 that establishes the connection between the MPH of the MCbiF and that of the Merge-Rips bifiltration constructed from the matrix of first-merge times, $D_{\theta,s}$.
Proposition 35.
@@
The proof of the proposition then follows from a simple extension of Proposition 36 in Schindler & Barahona (2025) to the 2-parameter case. To see that the 0-dimensional MPH of $\mathcal{L}$ and $\mathcal{M}$ are equivalent, note that both bifiltrations have the same 1-skeleton. Moreover, the 1-dimensional MPH is generally not equivalent because $\mathcal{L}$ is a Rips-based bifiltration and thus 2-determined, whereas $\mathcal{M}$ is not 2-determined.
If $\theta$ is strictly hierarchical, then $D_{\theta,s}$ fulfils the strong-triangle inequality, and thus the Rips-based bifiltration leads to a trivial 1-dimensional homology, see (Schindler & Barahona, 2025, Corollary 37). Hence, the 1-dimensional MPH of $\mathcal{L}$ is trivial. ∎
@@
Using Proposition 17, we prove the statement with the equivalent nerve-based MCbiF. Note that the graph $G:=P_{t_{2}|t_{1}}P_{t_{2}|t_{1}}^{T}$ has the same vertices and edges as the simplicial complex $\tilde{K}^{t_{1},t_{2}}$, which is 1-dimensional and thus also a graph according to Proposition 16. This shows that $\mathrm{HF}_{(}t_{1},t_{2})$ is given by the number of connected components in $G$. Furthermore, observe that $P_{t_{2}|t_{1}}^{T}\mathbf{1}=\mathbf{1}$, and that the resulting matrix $L$ is the Laplacian of the undirected graph $G$. Hence, $\dim(\ker L)$ is equal to the number of connected graph components (Chung, 1997), proving the statement. ∎
B.4 Proofs and Details for Section 6
@@
Appendix C Additional Examples
@@
Figure 8: Hilbert functions and optimised Sankey diagram for the sequence of partitions $\theta$ defined in Example 38.
@@
Let $X=\{x_{1},x_{2},x_{3},x_{4}\}$ for which we consider two different sequences of partitions $\theta(t)$ and $\eta(t)$ such that $\theta(1)=\eta(1)=\{\{x_{1},x_{2}\},\{x_{3}\},\{x_{4}\}\}$, $\theta(2)=\eta(2)=\{\{x_{1}\},\{x_{2},x_{3}\},\{x_{4}\}\}$ but $\theta(3)=\{\{x_{1},x_{3}\},\{x_{2}\},\{x_{4}\}\}\neq\theta(3)=\{\{x_{1}\},\{x_{2}\},\{x_{3},x_{4}\}\}$. See Fig. 9 for a Sankey diagram representation of the two sequences of partitions. Note that $\theta$ and $\eta$ only differ at scale $t=3$. However, this difference is crucial because a 1-conflict emerges in $\theta$ at scale $t=3$, whereas $\eta$ has only triangle 0-conflicts and no 1-conflict. Note that $\theta$ corresponds to the toy example in Fig. 1 with one additional isolated element.